OpenStudy (anonymous):

The set {0, 1} is closed under which operation? multiplication <-- my answer addition subtraction none of the above Question 2.2. Find A ∩ B A = {2, 3, 4, 5, 6} B = {3, 6, 9, 12, 15} {3, 6} <-- {2, 4, 5, 9, 12, 15} {2, 4, 5} {9, 12, 15}

OpenStudy (anonymous):

@mathstudent55

OpenStudy (amistre64):

looks fine to me

OpenStudy (mathstudent55):

Question 1. Multiplication and Addition. Not subtraction.

OpenStudy (mathstudent55):

Question 2.2 is correct.

OpenStudy (anonymous):

I can only pick one though :/ Which one would be better? Which expression is greater than |0.4|? |(0.4)| <-- ((0.4)) |(0.4)| |0.4| Question 4.4. Which number is one-fourth of the way between 3 and 11.4? 5.1 <-- 8.4 9.3 14.4

OpenStudy (mathstudent55):

4.4 is correct.

OpenStudy (anonymous):

Okay, is the one above that correct?

OpenStudy (mathstudent55):

In the first question multiplication and addition are equally correct, so picking multiplication is fine.

OpenStudy (anonymous):

Which expression is greater than |0.4|? |(0.4)| <-- ((0.4)) |(0.4)| |0.4|

OpenStudy (mathstudent55):

Does the double parentheses have a special meaning? I am not familiar with that.

OpenStudy (anonymous):

We did not discuss it in class so I am unsure. and I actually typed it wrong. Let me retype it.

OpenStudy (mathstudent55):

Also, choices A and C are the same.

OpenStudy (anonymous):

Which expression is greater than -|-0.4|? |-(-4)| -(-(-0.4)) -|-(-0.4)| - |-0.4|

OpenStudy (mathstudent55):

Oh, ok. This is much better.

OpenStudy (anonymous):

Yeah sorry. You need the negatives! Haha

OpenStudy (mathstudent55):

Any expression with 0.4 in it can only be 0.4 or -0.4, ok?

OpenStudy (anonymous):

Okay

OpenStudy (mathstudent55):

That means choices B, C, and D can be at most 0.4.

OpenStudy (mathstudent55):

Now let's look at choice A.

OpenStudy (anonymous):

Okay

OpenStudy (mathstudent55):

|-(-4)| = First, what is -(-4) = ?

OpenStudy (anonymous):

4

OpenStudy (mathstudent55):

Right, so you have |-(-4)| = |4| = What is absolute value of 4?

OpenStudy (anonymous):

4!

OpenStudy (mathstudent55):

Be careful. In math, the exclamation mark actually has a meaning. :-) 4! = 4 * 3 * 2 * 1 = 24 (Look up "factorial" in a math textbook.) I know what you mean. |4| = 4 Since the first expression has a value of 4, and all others can be at most 0.4, then the answer is A.

OpenStudy (anonymous):

Oh opps I didnt know that the exclamation mark meant anything! Okay awesome sauce! I just have two more that need checked :) Which Im sure are right I am very careful though. I like getting a good grade. :)

OpenStudy (mathstudent55):

Wait a minute. I just noticed that your question also changed when you retyped the problem.

OpenStudy (anonymous):

Hmm?

OpenStudy (mathstudent55):

Which expression is greater than -|-0.4|?

OpenStudy (anonymous):

Yes

OpenStudy (mathstudent55):

First let's see what the question is. -|-0.4| = -0.4

OpenStudy (mathstudent55):

That means any number in the choices that is larger than -0.4 is good.

OpenStudy (anonymous):

What is the value of (x + 6)(x 1) when x = 1? 6 1 0 <-- 5 Which property is shown? (5x 3) + (5x 9) = 5 x(3 + 9) Multiplicative Identity Distributive Property <-- Multiplicative Inverse Property Commutative Property of Multiplication

OpenStudy (anonymous):

Oh okay. Well Im sure A can be correct.

OpenStudy (mathstudent55):

Now let's see the choices again.

OpenStudy (mathstudent55):

Which expression is greater than -|-0.4| = -0.4? A |-(-4)| = 4 B -(-(-0.4)) = -0.4 C -|-(-0.4)| = -0.4 D -|-0.4| = -0.4 Only A is correct.

OpenStudy (anonymous):

Okay so if A were to be |-(-0.4)| it would still be correct?

OpenStudy (anonymous):

(I only have 15 mins left on my computer before it is my brothers turns.)

OpenStudy (anonymous):

@mathstudent55 ?

OpenStudy (mathstudent55):

(x + 6)(x 1) What does the second set of parentheses mean? I don't see an operation between x and 1.